# Asteria: Semantic-Aware Cross-Region Caching for Agentic LLM Tool Access

### v1

@@ 3.1 System Model and Goal
Deployment model. For large language model (LLM) agents that rely on external knowledge, frequent data retrieval from remote sources introduces significant latency and financial cost. This challenge is particularly pronounced when the serving LLM and knowledge base are deployed across regions rather than in the same on-premise cloud. In practice, agentic applications typically rely on MCP protocol [4] or other general RPC mechanisms [25] over wide-area networks to fetch the required data, amplifying these performance and cost penalties. To mitigate these overheads, we introduce Asteria, a novel cross-region caching system designed to store and serve frequently accessed knowledge from the perspective of agentic system efficiency.
Optimization Goal. The goal of Asteria is to reduce costly remote data accesses by reusing semantically equivalent knowledge. Instead of relying on literal text similarity or exact matching, Asteria leverages LLM-based semantic matching to identify and serve cached results for new queries with the same intent. This reduces query latency and significantly cuts API or tool usage costs, improving performance and operational efficiency in geo-distributed LLM agent deployments.
@@ 3.2 Design Opportunities and Challenges
Strawman design. As discussed in subsection 2.4, semantic prompt caches reuse model outputs based on embedding similarity. A natural strawman is to extend this idea: embed each cached retrieval query, place all keys in an ANN index, and on a new query, return the value attached to its nearest neighbor. ANN methods [31, 35] can search large embedding sets efficiently using graph structures or quantization. However, this naive design exposes key gaps. It treats similarity search as if it were a cache, yet offers no guarantee that top-ranked items are actually correct, fresh, or cost-efficient to reuse. These issues motivate the following design challenges.
Challenge 1: Ensuring accurate and valid semantic matches. Embeddings with ANN can reliably find and return textually similar queries, but textual or surface-level similarity [15] does not imply semantic equivalence. Two queries can be near in the vector space yet require different answers or reflect different intents. Correctness also depends on context and time. A match for a historical fact may remain valid, whereas a match for a fast-evolving topic can become stale. A practical system needs a validation mechanism that turns noisy similarity scores into trustworthy hit-or-miss decisions while accounting for both contextual fit and temporal validity.
@@
3.3 Asteria Architecture Overview
To address these three challenges, Asteria adopts a layered architecture shown in Figure 4. Each component is explicitly designed to tackle a subset of these challenges while working together with other components as an integrated system. The top tier is the user agentic application, where an LLM agent generates queries. The bottom tier is the Remote Data Service, the source of data, located in a remote region or cloud with a data retrieval latency around 300 ms-500 ms, following the common industrial practice [40]. Between these two lies the Asteria Engine, which provides a transparent caching interface. It consists of three primary internal components:
Data client. The data client serves as the transparent entry point into the Asteria system, which intercepts all outgoing requests generated by the agent before they can reach external services. This interception allows Asteria to seamlessly redirect the agent’s queries into its internal caching workflow without modifying the agent application.
Asteria cache engine. At the core of the engine is the retrieval pipeline, engineered to achieve high-precision semantic matching. This pipeline operates in two distinct stages to balance speed with accuracy. Initially, an ANN index [31] (Faiss [16] is used in our experiments.) based cache rapidly identifies a set of candidate data items based on vector similarity. These candidates are then passed to a lightweight semantic judger model (typically small with 1B parameters), which acts as a validation, being prompted to scrutinize each one for true contextual and semantic relevance to eliminate the false positives that plague naive semantic query caches.
@@ 3.3 Asteria Architecture Overview
Workflow The workflow begins when the data client transparently intercepts a structured request and its surrounding context generated by the LLM agent (➀-➁). Instead of immediately dispatching the request externally, it first queries the Asteria cache engine (➂). This query initiates a two-stage retrieval process designed for speed and accuracy. First, an ANN index rapidly surfaces a small set of candidate data items. Then, these candidates are scrutinized by the semantic judger (➆), ensuring accuracy. If the judger confirms a valid match, the cached result is returned instantly to the agent (➃). Otherwise, the request proceeds to the appropriate Remote Data Service (➄). The retrieved response is then returned to the agent and simultaneously stored in the cache as a new Semantic Element for future usage. Concurrently, the agent-aware cache manager observes the sequence of validated queries (➅), using this history to proactively prefetch data items (➇) and perform eviction, continuously optimizing the cache’s contents in the background.
4 Asteria’s Design
In this section, we present Asteria ’s architecture, highlighting how it transforms raw agent interactions into reusable knowledge units and orchestrates caching, retrieval, and scheduling to achieve fast, accurate, and cost-efficient performance.
@@ 4.1 Semantic Element
We begin with the Semantic Element (SE), the core building block and caching unit in Asteria. An SE organizes an agent’s query/action and the corresponding retrieval result so that later components can match, validate, and manage for reuse. As shown in Figure 5, an SE is a key–value pair: the agent’s query or tool action is the semantic key, and the retrieved information is the value. This structure leverages well-formed agentic outputs that wrap steps with tool tags. For example, the query “Who painted the Mona Lisa” within a <search> tag becomes the key, while the snippet in the <info> tag is the value. By reliably parsing these tagged blocks, Asteria encapsulates discrete interactions (web searches, API calls, database lookups) into coherent SEs for caching.
An SE also carries metadata essential for lifecycle decisions. First, an embedding model (e.g., Qwen3-0.6B [60]) converts the semantic key into a high-dimensional vector used as a semantic fingerprint and is used for future matching. Concurrently, a lightweight staticity score model (we reuse the semantic judge model in subsection 4.2) scores the staticity of the query–result pair on a 1–10 scale, indicating how fact-like and time-invariant the answer is. A higher value indicates a more fact-oriented response. For instance, the location of the Louvre Museum is stable (e.g., 10), whereas weather is ephemeral (lower). The system also records access frequency, retrieval latency, cost, and size (token length) per SE, quantifying the overhead of the original tool call. As detailed in subsection 4.3, these signals drive cost-aware eviction and prefetching.
4.2 Sine: Semantic Retrieval Index
We now describe how Asteria retrieves semantically appropriate matches efficiently via the Semantic Retrieval Index (Sine). Building on SEs, Asteria implements a high-throughput, dual-stage retrieval pipeline engineered to resolve the fundamental trade-off between retrieval speed and semantic precision. This system is a sophisticated interplay of multiple specialized models, and its core is a formal constrained optimization problem: to minimize end-to-end query latency while rigorously maintaining semantic accuracy.
To formalize this, we define the components and their associated latencies. The primary model is the agentic LLM, a powerful reasoning model that orchestrates tool use; its inference time is denoted as . The cache itself relies on two smaller, highly optimized models: a lightweight Embedding LLM for generating query vectors and a lightweight semantic judger model (LSM), a classifier that validates whether a cached result answers a new query and estimates its staticity. The combined latency of a full two-stage cache lookup/check is . On a cache miss, an external tool call adds .
@@ 4.2 Sine: Semantic Retrieval Index
Fine-grained validation. Selected candidates proceed to a high-precision semantic validation stage, where the LSM’s threshold, , is the primary lever for controlling accuracy (e.g., 0.9 in our experiment). The LSM evaluates if the cached result is a sufficient answer for the new query , producing a confidence score . A cache hit is confirmed only if . A higher enforces a stricter standard for equivalence, which increases the cache’s precision (fewer false positives) but can decrease the hit rate by rejecting marginally correct answers. A lower boosts the hit rate but at the direct risk of serving more incorrect results, thereby lowering precision.
@@
Optimization goal. The system minimizes expected latency subject to a target precision that is verified periodically offline. A hit incurs , while a miss incurs :
@@
Recalibration A fixed is brittle under workload drift or changing correctness requirements. Asteria therefore performs periodic offline recalibration to adjust the LSM decision boundary while keeping the agent’s latency-critical path unaffected. This process, detailed in algorithm 1, begins by creating a high-quality annotated dataset from a sample of recent queries (algorithm 1-1). For each sampled query, a Ground-Truth Evaluator provides the correct result, allowing the system to label the cache’s original answer as “correct” or “incorrect”. The objective is then to find a new threshold, , that aligns the cache’s performance on this dataset with a pre-defined metric, . This target precision represents the desired quality standard (e.g., 0.99 for 99% correctness). To find the new threshold, the system calculates a precision curve for the current LSM (algorithm 1-1) and identifies the value that satisfies (algorithm 1), which is then deployed to the live system.
@@
Crucially, this recalibration is highly cost-efficient: in our deployment it samples 5 recent queries per minute, adding negligible overhead. The annotated set can also fine-tune the LSM to better capture subtle semantic differences over time.
4.3 Building Cache Architecture Atop Sine
@@
Although the Sine index can identify semantically matched items, it is not by itself a cache. To bridge the gap, Asteria layers standard cache abstractions on top of the Sine, defining cache hit, cache eviction and cache prefetching and translate semantic matching signals into cache behavior.
@@
Semantic-aware cache hit. Unlike traditional caches where a hit is a simple key lookup, in our system a cached SE is only considered a “hit” after passing the full validation pipeline detailed in subsection 4.2. On a new query, ANN retrieves a small set of candidates from the cache. Then the LSM (semantic judger) validates equivalence. Only a valid match registers as a cache hit, and that increments an SE’s frequency count. This process effectively transforms a probabilistic similarity score into a definitive, binary hit/miss event, providing the stable, deterministic bedrock.
@@
LCFU eviction policy. With a reliable access signal established, we employ a tailored Least Cost-Efficient and Frequently Used (LCFU) eviction policy to manage the cache’s contents. The fundamental insight behind LCFU is that not all data are equally beneficial for overall financial and latency savings. Traditional LRU and LFU policies, which prioritize recency and frequency respectively, fail to account for the varying value of cached items. LCFU addresses this by assigning each SE a value_score that quantifies the benefit of retaining it.
As detailed in algorithm 2, this value_score is a composite metric derived from retrieval latency and cost, frequency, and staticity. Then the score is carefully normalized (algorithm 2). It adjusts the raw metrics for frequency and cost to ensure they always contribute positively and meaningfully. This prevents new items or those with low costs from being unfairly penalized, given the fact that the cost per request is less than 1, and taking its logarithm will return a negative value. This combined score is then normalized by the item’s size to provides a direct rationale: keep items that save the most time/money per byte.
For instance, data with a high Retrieval Cost and high access Frequency will naturally receive a higher value_score, making them less likely to be evicted. Conversely, ephemeral data with a lower Staticity will easily lead to lowering value_score, making it a more likely candidate for eviction even if its frequency is high. For one Stable data, with a large staticity and high retrieval cost, it will maintain a higher value_score even with less frequent access. This multi-attribute approach prevents transient but popular data from displacing enduring content, optimizing the cache for sustained performance and cost savings. When cache capacity is exceeded, items with the lowest value_score are evicted, preserving the most valuable content.
While a high value_score indicates valuable content, such entries can become stale if kept indefinitely. To prevent outdated information from persisting, Asteria integrates an aging mechanism using a user-defined Time-To-Live (TTL). Each cache entry is assigned a maximum lifespan, after which it is evicted regardless of its value_score. This ensures that even high-cost or frequently accessed items are periodically refreshed, maintaining the cache’s correctness while preserving the benefits of value-based retention.
@@
Predictive prefetching. To complement the reactive LCFU eviction policy, our framework employs a proactive history-based predictive prefetching to further reduce miss latency, leveraging the temporal correlations observed in agent workloads. This technique operates by learning query-to-query transition patterns from the stream of confirmed hits. Using a lightweight first-order Markov model, it calculates the probability that one query will follow another, as shown in Algorithm 3 (algorithm 3). A prefetch is triggered only if this probability exceeds a confidence threshold and the item is not already in the cache (algorithm 3). This initiates an asynchronous fetch to add the item to the cache (algorithm 3). A prefetched item enters the cache as a new SE with zero frequency. Consequently, its initial score is minimal. If the item is later requested and confirmed by the judger, its frequency is incremented, raising its value_score and making it a valuable, retained member of the cache. If it remains unused, its low score makes it a prime candidate for eviction by LCFU when space is required. This creates a low-risk, self-correcting system that minimizes cache pollution from speculative fetches.
@@ 4.4 Resource-Efficient Model Co-location
Although the two-stage retrieval computation is lightweight, the existing serving framework [63] adopts one model per GPU mode. To keep overall latency low and achieve high-precision without doubling hardware costs, Asteria co-locates the agent (e.g. 7B) and judger (e.g. 1B) efficiently on a single GPU.
@@
The central challenge is managing resource contention to protect the agent’s user-facing latency from the judger’s background processing. Our co-location architecture is built on a key insight: agent and judger workloads have fundamentally different priorities and resource profiles. Agent work is latency-critical, while judger work is a deferrable optimization. This is because a delayed validation does not block the user; at worst, it is treated as a cache miss, where performance for that single request degrades to the non-cached baseline of a full remote data call. Furthermore, the judger has a smaller memory demand since its prefill-only inference pattern (single token generation), results in a minimal and predictable KV cache footprint. Based on these principles, Asteria implements a two-level defense, shown in Figure 6, to ensure robust and efficient co-location.
@@
Coarse-grained asymmetric partitioning. The foundation of our co-location strategy is to exploit the fundamentally asymmetric nature of the agent and judger workloads. The judger’s workload profile is highly economical: unlike the agent which generates variable-length responses, the judger performs a classification task that yields a single token. This makes its primary consumer of GPU memory (i.e., the KV cache) minimal and highly predictable. Furthermore, its computational needs are lighter than the agent’s complex, multi-step reasoning. To capitalize on this, we use the CUDA Multi-Process Service (MPS) [48] to create a static, asymmetric compute partition, as shown in Figure 6. For instance, we allocate the dominant share of compute resources (e.g., 80%) to the agent and a smaller, sufficient share (e.g., 20%) to the judger. This coarse-grained partitioning optimizes for the common case, maximizing the performance of the primary agent LLM by dedicating the vast majority of resources to it.
@@
Fine-grained prioritization as a high-load guardrail. While static partitioning is effective for the common case, a second, dynamic layer of defense is required to handle periods of high contention and guarantee the agent’s low latency. This is the role of our fine-grained, priority-aware admission controller, which manages the unified dynamic memory pool (). This scheduler acts as a crucial guardrail by enforcing a strict prioritization policy. It services the agent queue () exhaustively, only considering a batch from the judger queue () when the agent queue is empty or lacks sufficient memory for the next dispatch. This ensures that the deferrable, internal work of the judger can never block the critical path of a user-facing agent task. This two-level defense—an asymmetric static partition for baseline efficiency and a dynamic priority scheduler for worst-case protection—is what enables robust, high-performance co-location.

### v2

@@ 3.1 System Model and Goal
Deployment model. For LLM-based agents that rely on external knowledge and tools, frequent data retrieval from remote sources introduces significant latency and financial cost. Geographic separation between inference and data sources is often inevitable: organizations calling external LLM APIs cannot control where providers deploy their models, and even with local inference, agents frequently invoke external tools (e.g., web search, databases, third-party APIs) hosted in remote regions. In practice, agentic applications typically rely on the MCP protocol [antropic2024mcp] or other general RPC [grpc2025] over wide-area networks to fetch required data, amplifying these performance and cost penalties. To mitigate these overheads, we introduce Cortex, a novel cross-region caching system designed to store and serve frequently accessed tool results from the perspective of agentic system efficiency.
Optimization Goal. The goal of Cortex is to reduce costly remote data accesses by reusing semantically equivalent knowledge. Instead of relying on literal text similarity or exact matching, Cortex leverages LLM-based semantic matching to identify and serve cached results for new queries with the same intent. This reduces query latency and significantly cuts API or tool usage costs, improving performance and operational efficiency in geo-distributed LLM agent deployments.
@@ 3.2 Design Opportunities and Challenges
Strawman design. As discussed in § 2.4, semantic prompt caches reuse model outputs based on embedding similarity. A natural strawman design is to extend this idea: embed each cached retrieval query, place all keys in an ANN index, and on a new query, return the value attached to its nearest neighbor. ANN methods [diskann, pq_ann] can search large embedding sets efficiently using graph structures or quantization. However, this naive design exposes key gaps. It treats similarity search as if it were a cache, yet it offers no guarantee that top-ranked items are actually correct, fresh, or cost-efficient to reuse. These issues motivate the following design challenges.
Challenge 1: Ensuring accurate and valid semantic matches. Embeddings with ANN can reliably find and return textually similar queries, but textual or surface-level similarity [doostmohammadi2023surfacebasedretrievalreducesperplexity] does not imply semantic equivalence. Two queries can be near in the vector space yet require different answers or reflect different intents. Correctness also depends on context and time. A match for a historical fact may remain valid, whereas a match for a fast-evolving topic can become stale. A practical system needs a validation mechanism that turns noisy similarity scores into trustworthy hit-or-miss decisions while accounting for both contextual fit and temporal validity.
@@
3.3 Cortex Architecture Overview
To address these three challenges, Cortex adopts a layered architecture shown in Figure 4. Each component is explicitly designed to tackle a subset of these challenges while working together with other components as an integrated system. The top tier is the user agentic application, where an LLM agent generates queries. The bottom tier is the Remote Data Service, the source of data, located in a remote region or cloud with a data retrieval latency around 300 ms-500 ms, following the common industrial practice [lin2024parrot]. This latency encompasses not only network RTT but also API processing and request queuing. Between these two lies the Cortex Engine, which provides a transparent caching interface. It consists of three primary internal components:
Data client. The data client serves as the transparent entry point into the Cortex system, which intercepts all outgoing requests generated by the agent before they can reach external services. This interception allows Cortex to seamlessly redirect the agent’s queries into its internal caching workflow without modifying the agent application.
Cortex cache engine. At the core of the engine is the retrieval pipeline, engineered to achieve high-precision semantic matching. This pipeline operates in two distinct stages to balance speed with accuracy. Initially, an ANN index [diskann] (Faiss [douze2024faiss] is used in our experiments.) based cache rapidly identifies a set of candidate data items based on vector similarity. These candidates are then passed to a lightweight semantic judge model (typically small with 1B parameters), which acts as a validation and is prompted to scrutinize each one for true contextual and semantic relevance to eliminate the false positives that plague naive semantic query caches.
@@ 3.3 Cortex Architecture Overview
Workflow The workflow begins when the data client transparently intercepts a structured request and its surrounding context generated by the LLM agent (➀-➁). Instead of immediately dispatching the request externally, it first queries the Cortex cache engine (➂). This query initiates a two-stage retrieval process designed for speed and accuracy. First, an ANN index rapidly surfaces a small set of candidate data items. Then, these candidates are scrutinized by the semantic judge (➆), ensuring accuracy. If the judge confirms a valid match, the cached result is immediately returned to the agent (➃). Otherwise, the request proceeds to the appropriate Remote Data Service (➄). The retrieved response is then returned to the agent and simultaneously stored in the cache as a new Semantic Element for future usage. Concurrently, the agent-aware cache manager observes the sequence of validated queries (➅), using this history to proactively prefetch data items (➇) and perform eviction, continuously optimizing the cache’s contents in the background.
4 Cortex’s Design
This section details Cortex’s architecture for transforming agent interactions into reusable knowledge units with efficient caching, retrieval, and scheduling.
@@ 4.1 Semantic Element
We begin with the Semantic Element (SE), which is Cortex’s core caching unit. An SE organizes an agent’s query and the retrieval result for matching, validation, and reuse. As shown in Figure 5, an SE is a key–value pair: the agent’s query or tool action is the semantic key, and the retrieved information is the value. This structure leverages well-formed agentic outputs that wrap steps with tool tags. For example, the query “Who painted the Mona Lisa” within a <search> tag becomes the key, while the snippet in the <info> tag is the value. By reliably parsing these tagged blocks, Cortex encapsulates discrete interactions (web searches, API calls, database lookups) into coherent SEs for caching.
An SE also carries metadata essential for lifecycle decisions. First, an embedding model (e.g., Qwen3-0.6B [qwen3_embedding_06b]) converts the semantic key into a high-dimensional vector used as a semantic fingerprint and is used for future matching. Concurrently, the semantic judge model (§ 4.2) estimates a staticity score that quantifies the expected validity duration of the query–result pair on a scale of 1–10: higher scores (9–10) indicate stable facts that are unlikely to change, while lower scores (1–3) indicate time-sensitive information. For instance, “Who painted the Mona Lisa?” is stable (10), “Who is the current US President?” is moderate (5), whereas “Today’s weather in Paris” is ephemeral (1). This score impacts the Time-To-Live assignment and the eviction priority in § 4.3.
4.2 Seri: Semantic Retrieval Index
Cortex retrieves semantic matches via via the Semantic Retrieval Index (Seri). Building on SEs, Cortex implements a high-throughput, dual-stage retrieval pipeline engineered to resolve the fundamental trade-off between retrieval speed and semantic precision. This system is a sophisticated interplay of multiple specialized models, and its core is a formal constrained optimization problem: to minimize end-to-end query latency while rigorously maintaining semantic accuracy.
To formalize this, we define the components and their associated latencies. The primary model is the agentic LLM, a powerful reasoning model that orchestrates tool use; its inference time is denoted as . The cache itself relies on two smaller, highly optimized models: a lightweight Embedding LLM for generating query vectors and a lightweight semantic judge model (LSM), a classifier that validates whether a cached result answers a new query and estimates its staticity. The combined latency of a full two-stage cache lookup/check is . On a cache miss, an external tool call adds .
@@ 4.2 Seri: Semantic Retrieval Index
Fine-grained validation. Selected candidates proceed to a high-precision semantic validation stage, where the LSM’s threshold, , is the primary lever for controlling accuracy (e.g., 0.9 in our experiment). The LSM evaluates whether the cached result is a sufficient answer for the new query , producing a confidence score . This guards against false positives from surface-level similarity, e.g., “Apple nutrition facts” vs. “Apple stock price” share keywords but require different answers. The LSM rejects such matches by reasoning about semantic equivalence. A cache hit is confirmed only if . A higher enforces a stricter standard for equivalence, which increases the cache’s precision (fewer false positives) but can decrease the hit rate by rejecting marginally correct answers. A lower boosts the hit rate but at the direct risk of serving more incorrect results, thereby lowering precision.
@@
Optimization goal. The system minimizes the expected latency subject to a target precision that is verified periodically offline. A hit incurs , while a miss incurs :
The formula reflects an accuracy-throughput trade-off: lower improves hit rate but risks incorrect results; higher preserves accuracy but increases remote call frequency.
@@
Recalibration A fixed is brittle under workload drift or changing correctness requirements. Cortex therefore performs periodic offline recalibration to adjust the LSM decision boundary while keeping the agent’s latency-critical path unaffected. This process, detailed in Algorithm 1, begins by creating a high-quality annotated dataset from a sample of recent queries (Algorithm 1-1). Ground truth is obtained by re-issuing sampled queries to the live tool and comparing tool responses against the cached results. The objective is then to find a new threshold, , that aligns the cache’s performance on this dataset with a pre-defined metric, . This target precision represents the desired quality standard (e.g., 0.99 for 99% correctness). To find the new threshold, the system calculates a precision curve for the current LSM (Algorithm 1-1) and finds the value that satisfies (Algorithm 1), which is then deployed to the live system. Recalibration is decoupled from the request-serving path and does not block agent requests.
@@
Crucially, this recalibration is highly cost-efficient: in our deployment, it samples 5 recent queries per minute, adding negligible overhead. The annotated set can also fine-tune the LSM to better capture subtle semantic differences over time.
4.3 Building Cache Architecture Atop Seri
@@
Although the Seri index can identify semantically matched items, it is not by itself a cache. To bridge the gap, Cortex layers standard cache abstractions on top of the Seri, defining cache hit, cache eviction and cache prefetching and translate semantic matching signals into cache behavior.
@@
Semantic-aware cache hit. Unlike traditional caches where a hit is a simple key lookup, in our system a cached SE is only considered a “hit” after passing the full validation pipeline detailed in § 4.2. On a new query, ANN retrieves a small set of candidates from the cache. Then the LSM (semantic judge) validates equivalence. Only a valid match registers as a cache hit, and that increments an SE’s frequency count. This process effectively transforms a probabilistic similarity score into a definitive, binary hit/miss event, providing a stable, deterministic bedrock for a caching system.
@@
LCFU eviction policy. With a reliable access signal established, we employ a tailored Least Cost-Efficient and Frequently Used (LCFU) eviction policy to manage the cache’s contents. The insight behind LCFU is that not all data are equally beneficial for cost and latency savings. Traditional LRU and LFU policies, which prioritize recency and frequency respectively, fail to account for the varying value of cached items. LCFU addresses this by assigning each SE a value_score that quantifies the benefit of retaining it.
As detailed in Algorithm 2, this value_score is a composite metric derived from retrieval latency and cost, frequency, and staticity. Each factor captures a distinct retention benefit: frequency reflects reuse likelihood, cost and latency reflect savings per hit, and staticity reflects expected validity duration. Then the score is carefully normalized (Algorithm 2). It adjusts the raw metrics for frequency and cost to ensure they always contribute positively and meaningfully. This prevents new items or those with low costs from being unfairly penalized, given the fact that the cost per request is less than 1, and taking its logarithm will return a negative value. This combined score is then normalized by the item’s size to provides a direct rationale: keep items that save the most time/money per byte. For instance, data with a high Retrieval Cost and high access Frequency will naturally receive a higher value_score, making them less likely to be evicted. Conversely, ephemeral data with a lower Staticity will easily lead to lowering value_score, making it a more likely candidate for eviction even if its frequency is high. For one Stable data, with a large staticity and high retrieval cost, it will maintain a higher value_score even with less frequent access. This multi-attribute approach prevents transient but popular data from displacing enduring content, optimizing the cache for sustained performance and cost savings. When cache capacity is exceeded, items with the lowest value_score are evicted, preserving the most valuable content. While a high value_score indicates valuable content, such entries can become stale if kept indefinitely. To prevent outdated information from persisting, Cortex integrates an aging mechanism using a user-defined TTL. Each cache entry is assigned a maximum lifespan, after which it is evicted regardless of its value_score. This ensures that even high-cost or frequently accessed items are periodically refreshed, maintaining the cache’s correctness while preserving the benefits of value-based retention.
@@
Predictive prefetching. To complement the reactive LCFU eviction policy, our framework employs proactive history-based prefetching to reduce miss latency. Using a lightweight first-order Markov model trained on confirmed cache hits, the system learns query-to-query transition patterns and calculates . When this probability exceeds a confidence threshold, the predicted query is asynchronously fetched and added to the cache. Prefetched items enter with zero frequency, giving them minimal value_score. If later requested and validated, their frequency increments and they become retained members. If unused, their low score makes them prime eviction candidates. This self-correcting design minimizes cache pollution from speculative fetches.
@@ 4.4 Resource-Efficient Model Co-location
Although the two-stage retrieval computation is lightweight, the existing serving framework [vllm_repo] adopts one model per GPU mode. To keep overall latency low and achieve high-precision without doubling hardware costs, Cortex co-locates the agent (e.g. 7B) and judge (e.g. 1B) efficiently on a single GPU. The challenge is protecting agent latency from judge processing. Our co-location architecture (Figure 6) is built on a key insight: agent and judge workloads have fundamentally different priorities and resource profiles. Agent work is latency-critical, while judge work is a deferrable optimization. This is because a delayed validation does not block the user; at worst, it is treated as a cache miss, where performance for that single request degrades to the non-cached baseline of a full remote data call. Furthermore, the judge has a smaller memory demand since its prefill-only inference pattern (single token generation), results in a minimal and predictable KV cache footprint. Atop these principles, Cortex builds a two-level defense, to ensure robust and efficient co-location.
@@
Coarse-grained asymmetric partitioning. The foundation of our co-location strategy is to exploit the fundamentally asymmetric nature of the agent and judge workloads. The judge’s workload profile is highly economical: unlike the agent which generates variable-length responses, the judge performs a classification task that yields a single token. This makes its primary consumer of GPU memory (i.e., the KV cache) minimal and highly predictable. Furthermore, its computational needs are lighter than the agent’s complex, multi-step reasoning. To capitalize on this, we use the CUDA Multi-Process Service (MPS) [cuda-mps] to create a static, asymmetric compute partition, as shown in Figure 6. For instance, we allocate the dominant share of compute resources (e.g., 80%) to the agent and a smaller, sufficient share (e.g., 20%) to the judge. This coarse-grained partitioning optimizes for the common case, maximizing the performance of the primary agent LLM by dedicating the vast majority of resources to it.
@@
Fine-grained prioritization as a high-load guardrail. While static partitioning is effective for the common case, a second, dynamic layer of defense is required to handle periods of high contention and guarantee the agent’s low latency. This is the role of our fine-grained, priority-aware admission controller, which manages the unified dynamic memory pool (). This scheduler acts as a crucial guardrail by enforcing a strict prioritization policy. It services the agent queue () exhaustively, only considering a batch from the judge queue () when the agent queue is empty or lacks sufficient memory for the next dispatch. This ensures that the deferrable, internal work of the judge can never block the critical path of a user-facing agent task. This two-level defense combines asymmetric partitioning for baseline efficiency with dynamic priority scheduling for worst-case protection, enabling robust, high-performance co-location.